### v2

# Automatically Labeling \$200B Life-Saving Datasets: A Large Clinical Trial Outcome Benchmark

Abstract
Background The global cost of drug discovery and development exceeds \$200 billion annually, with clinical trial outcomes playing a critical role in the regulatory approval of new drugs and impacting patient outcomes. Despite their significance, large-scale, high-quality clinical trial outcome data are not readily available to the public, limiting advances in trial outcome predictive modeling.
Methods We introduce the Clinical Trial Outcome (CTO) knowledge base, a fully reproducible, large-scale (around 125K drug and biologics trials), open-source of clinical trial information including large language model (LLM) interpretations of publications, matched trials over phases, sentiment analysis from news, stock prices of trial sponsors, and other trial-related metrics. From this knowledge base, we additionally performed manual annotation of a set of recent clinical trials from 2020-2024.
Results We evaluated the quality of our knowledge base by generating high-quality trial outcome labels that demonstrate strong agreement with previously published expert annotations, achieving an F1 score of 94 for Phase 3 trials and 91 across all phases. Additionally, we benchmarked a suite of standard machine learning models on our manually annotated set, highlighting the distribution shift of recent trials and the need for continuously updated labeling methods.
Conclusions By analyzing CTO’s performance on recent trials, we showed a need for recent, high-quality trial outcome labels. We release our knowledge base and labels to the public at https://chufangao.github.io/CTOD, which will also be regularly updated to support ongoing research in clinical trial outcomes, offering insights that could optimize the drug development process.
Introduction
A clinical trial is an indispensable step toward developing a new drug, involving human participants to test the drug’s efficacy and safety for treating target diseases. In 2022, drug discovery and development spending reached 244 billion dollars globally [1]. Among these, the clinical trial market reached \$44.3 billion in 2020 and is expected to grow to \$69.3 billion by 2028 [2]. Low efficacy, safety issues, and poor trial protocol design can lead to trial failures [3, 4, 5]. Eroom’s Law111reverse of “Moore’s Law” shows that the number of new FDA-approved drugs per billion US dollars of R&D spending has halved approximately every nine years since 1950, even with inflation adjustment [6]. Despite this decline in approval efficiency, the volume of drug and biologics trials continues to grow each year. However, obtaining high-quality, human-labeled trial data remains expensive and largely inaccessible due to the limited public availability of labels. Given these challenges, predicting trial outcomes in silico—using computational methods—could significantly enhance drug development efficiency.
The FDA does not release the clinical trial ID (NCTID) in documents of approved drug applications. Other public data sources, such as the ClinicalTrials.gov database with more than 500,000 historical trials [7, 8, 9], provide vital information for identifying trial recruitment and design information but does not provide a concrete label for trial outcome directly. The absence of a centralized, easily accessible knowledge base that consolidates clinical trial outcomes, drug approvals, and intervention phases poses a significant challenge for researchers and drug developers [10]. This fragmented information landscape hinders the development of accurate predictive models and can slow down the drug discovery and development process. Furthermore, the number of clinical trials conducted each year continues to rise, with their data dynamically evolving under the influence of various external factors, such as technological innovations and global events like the COVID-19 pandemic. This ever-changing landscape underscores the need for updated trial outcome labels to develop robust and high-performing clinical trial outcome prediction models.
Trial Outcome Definition: Clinical trial outcomes are multifaceted and have diverse implications. These outcomes can involve meeting the primary endpoint as defined in the study, advancing to the next phase of the trial, obtaining regulatory approval, impacting the financial outcome for the sponsor (either positively or negatively), and influencing patient outcomes such as adverse effects and trial dropouts. Our paper follows the previous conventions [11, 12, 3, 13, 14, 15] and defines the trial outcome as a binary indicator, showing whether the trial achieves its primary endpoints and can progress to the next stage of drug development. For Phase 1 and 2 trials, success may mean moving to the next phase, such as from Phase 1 to Phase 2, and from Phase 2 to Phase 3. In Phase 3, success is measured by regulatory approval.
Recently, initial efforts have been made to forecast specific aspects of clinical trials to enhance their outcomes. These efforts include employing electroencephalographic (EEG) measures to predict the effects of antidepressant treatments [16], optimizing drug toxicity predictions based on drug and target properties [17], and using phase II trial results to anticipate phase III trial outcomes [18]. Additionally, there is a growing interest in developing comprehensive methods for predicting trial outcomes. For example, predicting drug approvals for 15 disease groups by analyzing drug and clinical trial features using classical machine learning techniques [14], using multimodal drug structure and text information to predict outcomes based on a supervised set of data [11], an algorithm that computes the probability of technical success via asking experts a standardized questionnaire [19], and multimodal trial outcome prediction via omics, text, clinical trial design, and small molecule properties [15].
Despite these efforts, several limitations still impede the utility of existing trial outcome prediction models. Namely–the lack of transparency in the clinical trial labeling process. Fu et al.’s TOP [11] is one example of a publicly available expert-curated clinical trial dataset. However, it is also quite limited because it only contains around 12,000 human-labeled interventional small-molecule drug trials, most of which concluded prior to 2020. To date, we are unaware of any other large-scale, publicly available, open-source (fully reproducible) effort to compute trial outcome labels.
To address these limitations, we propose CTO, the first large-scale, publicly available, open-source, and fully reproducible dataset of clinical trial outcomes. We compiled a comprehensive trial knowledge base by integrating diverse data sources, including trial-related publications, news headlines, stock prices of trial sponsors, and other relevant trial metrics. From this knowledge base, we derive weakly supervised trial outcome labels using the CTO’s automated labeling framework, which aggregates labels from trial phase linkages, LLM interpretations on publications, sentiment analysis of news headlines, p-values, and trial status, among other indicators. We validate CTO ’s labeling accuracy against previously published trial outcomes, achieving an F1 score of 91 across all trial phases, closely aligned with human-annotated clinical trial outcomes. Additionally, aided by the CTO’s framework, we manually curated a set of approximately 10,000 recent trials completed between 2020 and 2024. To ensure reproducibility and keep pace with rapidly evolving trial data, we share monthly updates of our trial knowledge base and labels generated by CTO ’s automated labeling framework.
Figure 1: Overview of our CTO benchmark pipeline. (A) Trial selection criteria used to curate a targeted dataset from the CTTI, narrowing from an initial 513.4K trials to 125.8K trials in CTO to a final manually curated set of 11K trials. (B) Trial knowledge base creation pipeline, including extraction of linked PubMed abstracts, recent abstract searches, trial metrics, news articles, and stock prices. (C) Clinical trial outcome labeling process, integrating the CTO automated labeling framework assisted manual labeling to generate curated gold standard labels. (D) The manual curation process, our rule-based annotation criteria, followed by manual checking on challenging clinical trials.
Methods
Trial Selection Process
In our study, we focused primarily on clinical trials involving drugs and biologics. Starting with a dataset of approximately 513,485 clinical trials from the CTTI, we applied a series of selection criteria to create a refined subset as illustrated in Figure 1A. First, we selected only trials with interventions categorized as drugs or biologics, reducing the dataset to 209,124 trials. Next, we excluded active trials and those still recruiting to focus exclusively on completed studies, resulting in 145,464 trials. Subsequently, trials without phase information were removed, further narrowing the dataset to 125,840 trials, which represents the number of trials labeled using our CTO’s automated labeling framework. To construct a manually curated test set, we then selected trials completed after 2020, yielding a subset of approximately 30,900 recent trials. Importantly, we ensured that this set had no overlap with the TOP benchmark. Finally, we performed a sampling across different trial phases and conditions, producing a manually curated set of 11,012 trials.
Trial Knowledge Base Creation
Based on the CTTI data, we generate a trial knowledge base by aggregating and linking various data sources to the trials, as shown in Figure 1B. Our pipeline enables regular updates of the knowledge base given the updated CTTI data in a monthly frequency. The trial knowledge base extraction pipeline consists of four components: 1) PubMed abstracts extraction, 2) News extractions, 3) Trial metrics and 4) Stock price information extraction.
Pubmed abstract extraction: PubMed abstracts have been automatically linked to trials by the Clinical Trials Transformation Initiative (CTTI)[20, 21, 22] as well as other efforts [23]. Despite these efforts, we found that many recent abstracts were not linked to their corresponding trials. First, we extracted all linked PubMed abstracts for each trial through the NCBI API  [^2] https://ncbi.nlm.nih.gov/, and the statistics of the linked abstracts are given in the Supplementary. These abstracts can be categorized into 1) Background, 2) Derived, and 3) Results. Since we are interested in clinical trial outcomes, we only utilized abstracts in the Derived and Results categories. To further enrich the knowledge base, we performed regular searches on NCBI using NCT IDs to find newly published abstracts, linking any relevant ones to their associated trials. The initial count of drug or biologic trials with extracted PubMed abstracts linked from CTTI data was 49,277. Our additional search and extraction process linked 10,232 new trials with their recent PubMed abstracts. Ultimately, our knowledge base comprises 41,228 Results-type abstracts, 65,327 Derived-type abstracts, and 20,512 newly linked abstracts.
News extractions: In our initial analysis, we identified news sources as a valuable resource for determining the outcome of clinical trials, especially since the majority of trials in the CTTI dataset lack reported results. We extracted news articles related to completed drug and biologic trials by web scraping and linked them to the trials in our knowledge base. We sent requests to Google News for all completed drug and biologic trials (excluding phase 4), which accounted for around 85K trials. Due to rate limitations, we used an external API (SerpAPI) to streamline this process. Our findings revealed that most trials had no corresponding news coverage; only 3,688 trials were associated with any news articles.
Trial metrics: We extracted trial metrics from CTTI[24] as well as mining clinicaltrials.gov. Specifically, we extracted a total of 9 numerical metrics–number of patients, sponsors, sites, amendments, patient drops, deaths, serious adverse events, all adverse events, and the number of days the trial has been last updated after completion. We also extracted 3 binary metrics–whether results were reported, the completion/terminated status of the trial, and finally, whether the trial had a significant p-value. Detailed descriptions of each are in the supplementary.
Stock price information extraction: The stock price of a pharmaceutical or biotech company often reflects market expectations. If investors expect positive results, the stock may rise in anticipation of the trial’s completion. Conversely, if expectations are low or if previous trials have been unsuccessful, the stock may not perform as well. We utilized Yahoo finance [^3] https://pypi.org/project/yfinance/ to collect historical stock data for companies with publicly available tickers for completed trials. Details on the calculations involving stock price movements and their linkage to specific trials are provided in the supplementary.
CTO Automated Labeling Framework
We developed a clinical trial labeling pipeline that first assigns rule-based labels. First, trials with statuses indicating ‘terminated’, ‘withdrawn’, ‘suspended’, ‘no longer available’} were labeled as failures. Trials labeled as ’approved for marketing’ were marked as successes. For trials with ‘completed’ as status, we examined the reported results: if any reported p-value for a primary endpoint was $<$0.05, the trial was labeled as a success. Unlabeled trials are then automatically labeled. The main idea behind our automated labeling framework is to combine multiple sources of labels, which have varying levels of accuracy. We call these sources of labels “weakly supervised labeling functions” or LFs for short[25]. Following this, we apply a label model that leverages our sources to refine labels and produce denoised pseudo-labels. Each LF, such as detecting efficacy based on p-value thresholds, assigns labels or abstains when uncertain. Phase-specific thresholds are also applied to certain metrics, such as adverse events, to account for variations in participant numbers across trial phases.
From our trial knowledge base, we developed several sources of labels, including large language model (LLM) interpretations of PubMed abstracts, trial linkage across phases, sentiment analysis of news headlines, p-values, amendments, and other clinical trial metrics which were integrated into our clinical trial annotation pipeline as shown in Figure 1C. Further details on these labeling functions are given in the supplementary. Our pipeline is validated against several baselines, including a Majority Vote–which takes the majority vote over LFs, Data Programming–an unsupervised matrix completion approach that weighs LFs based on mutual agreement, and a supervised Random Forest model trained on LFs and a set of public labels (TOP).
Results
CTO Automated Labeling Achieves High Agreement with Previous Expert Labels
Table 1: Evaluation of automatically aggregated knowledge base data sources. Agreement of CTO with TOP (on TOP test split). $\text{CTO}_{MV}$ denotes Majority Vote, $\text{CTO}_{DP}$ denotes Data Programming [25], and $\text{CTO}_{RF}$ denotes Random Forest. Note that RF actively trained on the training human labels, whereas MV and DP are not. Bolded denotes best performance.
Phase Aggregation Method F1 $\kappa$
I $\text{CTO}_{MV}$ 0.726 0.490
$\text{CTO}_{DP}$ 0.870 0.700
$\text{CTO}_{RF}$ 0.913 0.790
II $\text{CTO}_{MV}$ 0.689 0.430
$\text{CTO}_{DP}$ 0.856 0.623
$\text{CTO}_{RF}$ 0.878 0.693
III $\text{CTO}_{MV}$ 0.904 0.606
$\text{CTO}_{DP}$ 0.921 0.582
$\text{CTO}_{RF}$ 0.941 0.710
All $\text{CTO}_{MV}$ 0.793 0.529
$\text{CTO}_{DP}$ 0.884 0.646
$\text{CTO}_{RF}$ 0.909 0.729
We evaluate our knowledge base and labeling framework by comparing the agreement between our automatically aggregated labels against the expert-annotated TOP test set. Table 1 summarizes the performance of various label aggregation methods across different trial phases. The label aggregation methods evaluated include Majority Vote (CTOMV), Data Programming (CTODP), and Random Forest-based aggregation (CTORF). Across all trial phases, the Random Forest-based approach consistently outperforms the other methods, achieving the highest scores in both F1 and Cohen’s kappa metrics. CTORF achieves an F1 score of $0.909$ and a Cohen’s kappa ($\kappa$) of $0.729$ when evaluated across all phases, indicating substantial agreement [26]. Also, the phase-specific analysis further underscores the robustness of CTORF. It achieves F1 scores of $0.913$, $0.878$, and $0.941$ in phases I, II, and III, respectively. Additionally, the F1 score of 0.909 is much higher than any previous SOTA trial outcome prediction model [27, 12].
Creating a Manual Curated Trial Outcome Benchmark from CTO
Since automated labeling is never a replacement for humans, we manually annotated over 2,500 challenging cases using our knowledge base to create a high-confidence gold standard set for evaluation. Our process of creating this benchmark is shown in Figure 1D. We follow the same rule-based termination and p-value$<$0.05 as in the automated labeling step to remove the easy cases. For the rest of the unlabeled trials, we utilized 3 clinical trial experts to label the unlabeled trials by assessing related publications, news articles, and other reported outcomes to determine the trial outcome. In this benchmark, we evaluated the labels generated by the CTO automated labeling framework against human labels. The CTO framework demonstrated strong overall performance, achieving an F1 of 97.1. Notably, even in hard and challenging cases (after p-value and status filtering), where accurate labeling is more complex, CTO maintained an F1 of 72.7.
Figure 2: Overview of our manually curated trial outcome benchmark. (A) Bar plot illustrating the top 10 conditions in the benchmark, categorized by MeSH ancestor terms, with associated label counts. (B) Distribution of trial phases and their respective outcome labels. (C) Annual distribution of completed trials in the set from 2020-2024 and their label counts.
Updated Trial Outcome Labels are Highly Valuable
Figure 3: (A) Year-by-year comparison of trial outcome prediction performance for baseline models trained on the TOP dataset, the constantly updated CTO dataset, and our manually curated set, evaluated on trials completed from 2021 to 2024. The TOP dataset is fully utilized for training, whereas for the CTO and manually curated sets, only trials completed prior to each test year are used for training. (B) Top 10 Condition-wise performance of baselines on manual labels, with models trained on data $<$ 2022 and tested on $\geq$ 2022 using TOP, CTO, and manually curated set. (C) Similar to B, Condition-wise performance on the next 10 (Top 11-20 conditions)
Given that we have manual labels for recent trials, we conducted experiments comparing the impact of our constantly updated CTO dataset against the previously publicly available fixed benchmark TOP dataset. We trained standard baseline trial outcome prediction models on the entire TOP dataset (train + valid + test) and assessed their year-by-year performance on trials completed between 2021 and 2024 in our manually curated dataset. For the CTO dataset, the baselines were trained using only trials completed prior to each test year, with training labels generated through our automated labeling framework. Similarly, we trained the baselines on the manually curated dataset using human-labeled outcomes for trials completed before each test year. The phase-wise performance comparison of these baseline models across the different training datasets is presented in Figure 3A.
Baseline models trained using labels from the CTO dataset consistently outperform their counterparts trained on the TOP dataset across all years and trial phases. The CTO dataset significantly enhances model performance compared to the TOP dataset, underscoring the importance of constantly updated labels. This improvement is largely attributable to distribution shifts caused by evolving trial practices, therapeutic areas, and advancements in trial designs over time.
This shows that CTO provides a scalable solution for labeling recent trial data, overcoming the infeasibility and time demands of obtaining new human labels. By achieving high agreement with human labels, CTO enables models to obtain close to human-level performance through leveraging large datasets to adapt to the evolving landscape of clinical trials.
Automated Labels for Condition-Specific Outcome Prediction
Figure 3B and C present the performance of baseline models trained using our human-labeled data across the top 20 conditions in the dataset (3B shows conditions 1-10 and 3C shows conditions 11-20). Here, we train the baselines on trials completed before 2022 over TOP, CTO, and manually curated. The baselines are tested on the manually curated trials completed after 2022. The performance of the baseline models trained on CTO closely follows the models trained on our human labels. This demonstrates that the labels produced by CTO align closely with manual annotations despite being generated independently through an automated process. Moreover, these CTO-based models consistently outperform those trained on the TOP dataset, demonstrating the value of using a more current dataset. Furthermore, CTO facilitates the development of condition-specific clinical trial outcome prediction models by rapidly generating automated labels for condition-specific datasets, significantly reducing the time and effort involved in manual labeling, with CTO-trained baselines matching or even exceeding baselines trained on manually-curated in some cases.
Discussion
Figure 4: Overview of the CTO dataset and labeling functions. (A)Trial distribution by completed year in CTO, which shows a steady increase over the years. (B) Frequency of top 30 conditions in the CTO dataset, categorized by MeSH ancestor terms. (C) Heatmap shows the agreement between the important labeling functions in CTO. (D) Distribution of trial phases and their counts in CTO.
The CTO dataset provides a comprehensive view of clinical trial data, capturing key information across multiple dimensions for around  125K drug and biologics trials. Figure 4A, B, and D provides an overview of the dataset’s key statistics, highlighting its scale and diversity. Figure 4A illustrates the annual distribution of completed clinical trials, and it shows a steady increase over the years, reflecting the growing scale of global clinical research. In Figure 4B, we show the frequency distribution of the top 30 medical conditions within the dataset, where Neoplasms are the most represented condition. The phase distribution of trials in the CTO dataset is shown in Figure 4D. Note that we ensured that our set of manual labels accurately reflects the overall clinical trial distribution as shown in Figure 2A, B and C.
How Well the Labeling Functions Agree?
To assess the agreement between the labeling functions, we calculated the percentage of agreement for trials that fall within the intersections of their coverage, as shown in Figure 4C. The heatmap highlights the agreement scores only for the important labeling functions in the dataset. Each cell in the heatmap represents the agreement score between a pair of labeling functions. It is important to note that each labeling function does not cover all trials with weak labels. Therefore, while calculating the agreement score, we only considered the trials common to both labeling functions. For example, there are no common trials with weak labels between the news headlines and trial status labeling functions (represented as a black cell). This is because the trial status labeling function assigns a value of -1 (indicating an inability to decide the trial outcome) to trials with a ‘completed’ status, while news extraction is limited to ‘completed’ trials to optimize resource usage associated with SerpAPI.
From the heatmap in Figure 2F, several key patterns in labeling function agreement can be observed. Notably, there is a high level of agreement between news headlines and labels derived from p-values, amendments, and serious adverse events. Additionally, we observe a strong agreement between trial linkage and trial status, suggesting that trial linkage effectively identifies failure trials, as most trial status-based labels pertain to trials categorized as terminated,’ withdrawn,’ or ‘suspended.’
Another noteworthy observation is the low agreement between p-values and trial status. This discrepancy arises because most trial status-based failure labels correspond to trials marked as ‘terminated,’ ‘withdrawn,’ or ‘suspended,’ often due to sponsor or business decisions rather than scientific outcomes. Interestingly, many of these trials report statistically significant p-values ($<0.05$) for some of their primary endpoints but are stopped for reasons unrelated to efficacy. For instance, our analysis identified approximately 1,200 trials marked as ‘terminated’ that reported primary outcome p-values below 0.05. Specific examples include NCT04364763, which was terminated due to enrollment feasibility issues, and NCT01374451, which was stopped for failing to meet the primary endpoint of progression-free survival. These cases highlight the complex interplay between trial outcomes and external factors.
Conclusion
We successfully created the first large-scale, publicly available knowledge base of clinical trial outcomes that automatically produces useful outcome labels. We additionally release a set of manually annotated outcome labels using human experts. Since we are the first to publicly release our trial annotation pipeline along with a corresponding knowledge base, we hope to democratize trial outcome prediction, enabling faster, more transparent trial design. Our dataset of recent trials, weak labels, and knowledge base is available at https://chufangao.github.io/CTOD/.
Acknowledgement
This work was supported by NSF award SCH-2205289, SCH-2014438, and IIS-2034479.
A.1    Motivation
  • For what purpose was the dataset created?
    We created CTO to democratize clinical trial outcome prediction, which was previously only available to industry-sponsored researchers. Furthermore, we attempt to expand the previous labeling efforts (which were primarily focused on small drug interventions) and predict labels on all trials.
  • Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
    The authors of this paper.
  • Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
    This work was supported by NSF award SCH-2205289, SCH-2014438, and IIS-2034479.
A.2    Composition
  • What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)?
    CTO contains molecule SMILES strings, eligibility criteria, ICD Codes, drug names, diseases, study status, phase information, and our automatically created labels. Furthermore, we also have additional QA tasks as extracted by GPT 3.5 Turbo 0125, news articles mined for the top 1000 industry sponsors, and stock prices.
  • How many instances are there in total (of each type, if appropriate)?
    There is a total of 479,761 trials, each with multiple types of predicted labels via random forest and data programming, as well as each phase-optimized threshold. For small-molecule drug interventions, we also have SMILES and ICD Codes. We release all such labels to not limit any downstream use. Furthermore, there are a total of 1,115,017 news articles we extracted and 105,570 trials with corresponding QA pairs.
  • Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
    We provide all possible trial labels up to the beginning of May 2024.
  • What data does each instance consist of?
    CTO contains molecule eligibility criteria, drug names, diseases, study status, phase information, and our automatically created labels.
  • Is there a label or target associated with each instance?
    The automatically predicted label is provided for each question.
  • Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
    No.
  • Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?
    No.
  • Are there recommended data splits (e.g., training, development/validation, testing)?
    Yes, please see Table 3 for a split of training trials with completion date $<2022$ tested on trials with completion date $\geq 2022$.
  • Are there any errors, sources of noise, or redundancies in the dataset?
    Automatically created labels inherently come with an element of noise. However, our high agreement with TOP’s human labels (up to 0.91 F1), implies that our labels are of high quality.
  • Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
    CTO depends on multiple open source datasets.
    1. 1.
    2. 2.
    3. 3.
  • Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)?
    No. We obtained all data sources via open-source methods.
  • Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
    No.
  • Does the dataset relate to people?
    Yes.
  • Does the dataset identify any subpopulations (e.g., by age, gender)?
    Yes, but only in the eligibility criteria of the trials, which are public.
  • Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals race or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)?
    No. There is no reference to individuals.
A.3    Collection process
  • How was the data associated with each instance acquired?
    We automatically mined each LLM prediction, trial linkage, news headline, and stock price as an overview in Section CTO Overview.
  • What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)?
    We mainly used Google Sheets and Python to collect, process, and label the data. In addition, we used OpenAI’s ChatGPT (GPT-3.5-turbo 0125) to generate QA and GPT outcome predictions.
  • If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
    Some data splitting was done according to previous data splits from TOP https://github.com/futianfan/clinical-trial-outcome-prediction. Additionally, we also split the data according to the following dates: $(-\infty,2018,2020,\infty)$
  • Who was involved in the data collection process (e.g., students, crowd workers, contractors) and how were they compensated (e.g., how much were crowd workers paid)?
    The authors of the paper collected and processed the data.
  • Over what timeframe was the data collected?
    We collected the data between December 2023 and April 2024.
  • Were any ethical review processes conducted (e.g., by an institutional review board)?
    N/A.
  • Does the dataset relate to people?
    Yes.
  • Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
    N/A.
  • Were the individuals in question notified about the data collection?
    N/A.
  • Did the individuals in question consent to the collection and use of their data?
    N/A.
  • If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses?
    N/A.
  • Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted?
    The dataset does not have individual-specific information.
A.4    Preprocessing/cleaning/labeling
  • Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?
    N/A.
  • Was the “raw” data saved in addition to the preprocess/cleaned/labeled data (e.g., to support unanticipated future uses)?
    N/A.
  • Is the software that was used to preprocess/clean/label the data available?
    Preprocessing, cleaning, and labeling are done via Google Sheets and Python.
A.5    Uses
  • Has the dataset been used for any tasks already?
    No.
  • Is there a repository that links to any or all papers or systems that use the dataset?
    No.
  • What (other) tasks could the dataset be used for?
    Our dataset is designed to promote research primarily in clinical trial outcome prediction. The dataset can also be used for stock price/trend prediction, question answering, etc.
  • Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
    N/A.
  • Are there tasks for which the dataset should not be used?
    N/A.
A.6    Distribution
  • Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
    No.
  • How will the dataset be distributed?
    Since clinical trial data is frequently updated, we provide the code for generating our CTO dataset at https://github.com/chufangao/CTOD. The current version of the dataset can be accessed at https://zenodo.org/doi/10.5281/zenodo.11535960.
  • Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
    The dataset is released under MIT License.
  • Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
    No.
  • Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
    No.
A.7    Maintenance
  • Who will be supporting/hosting/maintaining the dataset?
    The authors of this paper.
  • How can the owner/curator/manager of the dataset be contacted(e.g., email address)?
    Contact the corresponding authors (chufan2@illinois.edu & jp65@illinois.edu & trishad2@illinois.edu & jimeng@illinois.edu).
  • Is there an erratum?
    No.
  • Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
    If any corrections are needed, we plan to upload an updated version of the dataset along with detailed explanations of the changes.
  • If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were the individuals in question told that their data would be retained for a fixed period of time and then deleted)?
    N/A
  • Will older versions of the dataset continue to be supported/hosted/maintained?
    Primarily, we aim to keep only the latest version of the dataset. However, in specific cases like major updates to the dataset or the necessity to validate previous research with older versions, we will exceptionally retain past versions of the dataset for up to one year.
  • If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
    Contact the authors of this paper or raise a github issue.